# Linear Combinations of Logarithms of $L$-functions over Function Fields at Microscopic Shifts and Beyond

### v2

@@ 3.3. Dirichlet Polynomial Approximation for $\log{L(s,\mathcal{F})}$
By [15, (4.7)] for $s=\sigma+it$ with $\operatorname{Re}(s)\geq\frac{1}{2}$ (the case $\mathcal{F}=E\otimes\chi_{D}$ follows the same path, but it requires further information about zeros),
@@ 4.1. Microscopic Regime
Following the proof in [15, Proposition 6.1] and using (3.13), we obtain
@@
for a subset $\mathcal{H}^{\prime}_{n}\subset\mathcal{H}_{n}$ with $|\mathcal{H}_{n}\setminus\mathcal{H}^{\prime}_{n}|=o(1)$. This follows from the proof of [15, Proposition 6.1] together with Lemma 3.11. ∎
@@ 6.1. Correlations of the Real Parts of Dirichlet Polynomials
Take $\sigma_{0}=\frac{1}{2}+\frac{c}{X}$. We begin by using the Proposition 3.9 to write $\log|\mathfrak{L}_{\vec{a},\vec{t}}\left(s,\chi_{D}\right)|$ as a linear combination of $\mathcal{R}_{X,t_{j},D}$’s and then apply Proposition 6.1. Ve view $\log|\mathfrak{L}_{\vec{a},\vec{t}}\left(s,\chi_{D}\right)|$ as sum of two random variables, say $Y_{D}+Z_{D}$, where $Y_{D}=a_{1}\mathcal{R}_{X,t_{1},D}+\cdots+a_{k}\mathcal{R}_{X,t_{k},D}$ and $Z_{D}$ is an error term. The hypothesis $g\left(\sigma_{0}-1/2\right)=o(\sqrt{\log n})$ imply that $\log X=\log n+O(\log_{2}n)$, which satisfies the hypothesis of Proposition 6.1. Thus, by Proposition 6.1 the proof of the theorem is complete once we are able to show that (see [19, Lemma 2.9])
@@
This directly follows from combining Theorem 1.1 and Corollary 4.2. ∎
@@ 7. Proofs of Unconditional Results on the Critical Circle
As an application of [10, Theorem 2] and (7.1), we have the following corollary.
@@
Let $K=K_{Sp,2\pi}$ be the reproducing kernel of the Hilbert space associated to $\mathfrak{H}_{Sp,2\pi}$ given by [10, Theorem 6, case (ii)]. Then
@@
Its maximum value is its limit as $\alpha\to 0^{+}$, and $r(0)=0.9427\ldots$, as in [7]. The limit of $r(\alpha)$ is $\frac{1}{2}$ as $\alpha\to\infty$ ((see [10, Figure 5] for more details)).
@@
From [11, Corollary 10.1], one can obtain an asymptotic formula for the one-level density as
@@
As an application of [10, Theorem 2] and (7.3), we have the following corollary.
@@
Let $k=K_{\text{SO(even)},2\pi}$ be the reproducing kernel of the Hilbert space associated to $\mathfrak{H}_{\text{SO(even)},2\pi}$, determined by Case $(ii)$ of [10, Theorem 5]. Then
@@
Then the maximum value of $r_{E}(\alpha)$ is $1/4$ which is attained at $\alpha=0$, and its limit as $\alpha\to\infty$ is 0. We refer the reader to [10, Figure 3] for more details.
@@
where $\theta_{j}:=\theta_{j,D}(\mathcal{F})$’s are as defined in (4.2). Note that the summands in the second term are all nonnegative which leads to the negativity of the second term of the above expression. Therefore, we deduce that
@@ Acknowledgments
The first author was partially supported by a postdoctoral fellowship of the Pacific Institute for the Mathematical Sciences. The second author acknowledges financial support from the Australian Research Council Grant DP230100534 and the Max Planck Institute for Mathematics, Bonn.

### v3

@@ 3.3. Dirichlet Polynomial Approximation for $\log{L(s,\mathcal{F})}$
By [22, (4.7)] for $s=\sigma+it$ with $\operatorname{Re}(s)\geq\frac{1}{2}$ (the case $\mathcal{F}=E\otimes\chi_{D}$ follows the same path, but it requires further information about zeros),
@@ 4.1. Microscopic Regime
Following the proof in [22, Proposition 6.1] and using (3.13), we obtain
@@
for a subset $\mathcal{H}^{\prime}_{n}\subset\mathcal{H}_{n}$ with $|\mathcal{H}_{n}\setminus\mathcal{H}^{\prime}_{n}|=o(1)$. This follows from the proof of [22, Proposition 6.1] together with Lemma 3.11. ∎
@@ 6.1. Correlations of the Real Parts of Dirichlet Polynomials
Take $\sigma_{0}=\frac{1}{2}+\frac{c}{X}$. We begin by using the Proposition 3.9 to write $\log|\mathfrak{L}_{\vec{a},\vec{t}}\left(s,\chi_{D}\right)|$ as a linear combination of $\mathcal{R}_{X,t_{j},D}$’s and then apply Proposition 6.1. Ve view $\log|\mathfrak{L}_{\vec{a},\vec{t}}\left(s,\chi_{D}\right)|$ as sum of two random variables, say $Y_{D}+Z_{D}$, where $Y_{D}=a_{1}\mathcal{R}_{X,t_{1},D}+\cdots+a_{k}\mathcal{R}_{X,t_{k},D}$ and $Z_{D}$ is an error term. The hypothesis $g\left(\sigma_{0}-1/2\right)=o(\sqrt{\log n})$ imply that $\log X=\log n+O(\log_{2}n)$, which satisfies the hypothesis of Proposition 6.1. Thus, by Proposition 6.1 the proof of the theorem is complete once we are able to show that (see [28, Lemma 2.9])
@@
This directly follows from combining Theorem 1.1, Corollary 4.2, and Lemma 2.9 of [28]. ∎
@@ 7. Proofs of Unconditional Results on the Critical Circle
As an application of [15, Theorem 2] and (7.1), we have the following corollary.
@@
Let $K=K_{Sp,2\pi}$ be the reproducing kernel of the Hilbert space associated to $\mathfrak{H}_{Sp,2\pi}$ given by [15, Theorem 6, case (ii)]. Then
@@
Its maximum value is its limit as $\alpha\to 0^{+}$, and $r(0)=0.9427\ldots$, as in [12]. The limit of $r(\alpha)$ is $\frac{1}{2}$ as $\alpha\to\infty$ ((see [15, Figure 5] for more details)).
@@
From [18, Corollary 10.1], one can obtain an asymptotic formula for the one-level density as
@@
As an application of [15, Theorem 2] and (7.3), we have the following corollary.
@@
Let $k=K_{\text{SO(even)},2\pi}$ be the reproducing kernel of the Hilbert space associated to $\mathfrak{H}_{\text{SO(even)},2\pi}$, determined by Case $(ii)$ of [15, Theorem 5]. Then
@@
Then the maximum value of $r_{E}(\alpha)$ is $1/4$ which is attained at $\alpha=0$, and its limit as $\alpha\to\infty$ is 0. We refer the reader to [15, Figure 3] for more details.
@@
where $\theta_{j}:=\theta_{j,D}(\mathcal{F})$’s are as defined in (4.2). Note that the summands in the second term are all nonnegative which leads to the negativity of the second term of the above expression. Therefore,
@@ Acknowledgments
The first author was partially supported by a postdoctoral fellowship of the Pacific Institute for the Mathematical Sciences. The second author acknowledges financial support from the Australian Research Council Grant DP230100534 and the Max Planck Institute for Mathematics, Bonn. The authors all thank Jon P. Keating for his quick comments on this paper and also Paul Bourgade for suggesting a study of the FHK conjecture as in Section 1.3.3.